OpenStudy (anonymous):

Gasoline costs $2.99 per gallon. if the price per gallon increases an average of 6% per month, write a function that models the exponential growth of the pricing. a) f(x)=1.06(2.99)^x b) f(x)=2.99(1.06)^x c) f(x)=[1.06(2.99)]^x d) f(x)= 2.99/(1.06^x)

OpenStudy (kropot72):

Let x = number of months from the starting month. Each month the price is multiplied by 1.06, giving the 6% increase. So after the first month the price becomes \[2.99\times 1.06\] and after the second month the price becomes \[2.99\times 1.06\times 1.06\] and for the third month the second month price is multiplied by 1.06 giving \[2.99\times 1.06\times 1.06\times1.06\] Can you see the correct choice now?

OpenStudy (anonymous):

yes, i think it is B

OpenStudy (kropot72):

Yes, B is the correct choice.

OpenStudy (anonymous):

thanks for your help! :)

OpenStudy (kropot72):

You're welcome :)